# Instability of Legendrian knottedness, and non-regular Lagrangian concordances of knots

### v1

@@ 3.1. Decomposable fillings
A particularly nice class of Lagrangian cobordisms in ×$\mathbb{R}\times S^{3}$ are the decomposable ones, which were defined by Chantraine in [4]. By definition, a Lagrangian cobordism is decomposable if it consists of a concatenation of the following simple types of Lagrangian cobordisms:
@@
Figure 9. Ambient Legendrian surgery: The Legendrian is deformed in a neighborhood of the surgery arc (shown on top) to yield the Legendrian $\Lambda_{\eta}$ on the bottom. There is a standard Lagrangian 1-handle cobordism with concave boundary the Legendrian $\Lambda$ before the surgery, and convex boundary the Legendrian $\Lambda_{\eta}$ produced by the surgery.
@@
Decomposable Lagrangian cobordisms where shown to be regular in the work of Conway, Etnyre and Tosun [6]. It is still unknown whether the converse is true or not. The proof in [6] moreover gives an explicit recipe for finding the attaching spheres on the complement of the Legendrian. We proceed to explain this recipe in the case of a decomposable Lagrangian filling.
The 0-handles: For each zero-handle we consider $(\mathbb{C}^{2},-d^{c}\rho)$ with the standard filling $\mathfrak{Re}\mathbb{C}^{2}\subset\mathbb{C}^{2}$ of the standard Legendrian unknot $\partial_{\infty}\mathfrak{Re}\mathbb{C}^{2}\subset S^{3}$ of $\tt{tb}=-1$. This filling is clearly is tangent to the radial Liouville vector field, where $\rho(\mathbf{z})=\frac{\|\mathbf{z}\|^{2}}{4}.$ In the general case, when the Lagrangian cobordism has several 0-handles, one has to consider a Weinstein connected sum of such Weinstein four-balls. Since this Weinstein connected sum again produces the standard sphere, and since these handles can be determined from the number of zero-handles, we will omit these 1-handles from the diagram.
The 1-handles: We add a pair of cancelling Weinstein one and two-handles just above the region where the cusp-connected sum takes place; this is move K–I shown in Figure 4. The resulting diagram after having performed the cusp-connected sum is shown at the top of Figure 5. Before the cusp-connected sum, the two facing cusp-edges are now located inside the Weinstein 1-handle. In other words, the cusp-connected sum now takes place entirely inside the one-handle. It can readily be seen that the corresponding standard Lagrangian one-handle in this position is tangent to the standard Liouville structure on the Weinstein one-handle. In particular, the critical point of the Weinstein one-handle coincides with the critical point of the standard Lagrangian one-handle. See Proposition 3.2 for the general statement in arbitrary dimensions.
@@ 3.2. The case of a general ambient Legendrian surgery
of the zero-section in the above standard neighbourhood, while the Legendrian $\Lambda$ is identified with the Lagrangian co-normal bundle $N^{*}_{(\partial\eta)\times D^{n-k-1}}=N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$ of $(\partial\eta)\times D^{n-k-1}$, lifted to have vanishing $[-\epsilon,\epsilon]_{z}$-coordinate.
The choice of conformal symplectic normal bundle induces the identification of the $D^{*}D^{n-k-1}$-factor in the above neighbourhood. Since $N^{*}_{(\partial\eta)\times D^{n-k-1}}$ is a trivial $\mathbb{R}$-fibration over $(\partial\eta)\times D^{n-k-1}$, the compatibility of the framing of $\partial\eta$ and the above conformal symplectic normal bundle implies that the $\mathbb{R}$-fibre of the conormal bundle together with the standard coordinates on the $0_{D^{n-k-1}}$-factor together yield the framing of $\partial\eta\subset\Lambda$.
Passing to the symplectisation, we can find a neighbourhood of $[0,1]\times\eta\subset\mathbb{R}\times Y$ in the symplectisation of the form
@@
in which the cone $[0,1]\times\eta$ is identified with $0_{[0,1]}\times 0_{\eta}\times\{0\}$ and the cone $[0,1]\times\Lambda$ is identified with $0_{[0,1]}\times N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$. The standard Lagrangian $k+1$-handle is constructed to be a cylinder over $\Lambda$ outside the above neighbourhood of the symplectisation, while the intersection with the above neighbourhood is an explicitly constructed neighbourhood of the core-disc; see [9, Section 4.2.2]. We will only need a description of $L_{\eta}$ in a neighbourhood of the zero-section
@@
Namely, in a small neighbourhood of this zero-section, the standard handle $L_{\eta}$ is given by the Lagrangian co-normal bundle of a $k+1$-dimensional disc in $\Gamma\times\{0\}\subset[0,1]\times\eta\times D^{n-k-1}$ where $\Gamma\subset[0,1]\times\eta$ is a smooth codimension one embedding of a disc given as the union of $[0,1/2]\times\partial\eta$ and a suitable graph of a function $\eta\to[1/2,1]$. In particular, this means that $L_{\eta}$ intersects the zero-section of this standard neighbourhood in a representative $\overline{C}_{\eta}\subset\overline{L}_{\eta}$ of the $k+1$-dimensional core-disc of the handle. Further, we note that the core disc $\overline{C}_{\eta}$ is the boundary of an isotropic $k+2$-disc $\overline{D}_{\eta}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$ given by the sub-level set of the above graph. Here $\overline{D}_{\eta}$ is homeomorphic to a disc and has boundary with corners; one boundary stratum is the disc $\overline{C}_{\eta}$, while the other stratum is $\{0\}\times 0_{\eta}\times\{0\}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$. By
@@
    $L_{\eta}$ is contained in the $k+1$-handle and tangent to the Liouville flow, where the critical point of the Liouville flow of the handle coincides with the critical point of the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate;
@@
endowed with a Liouville form for which $\{(0,0)\}$ is a unique critical point of index $i$, whose descending manifold (i.e. the skeleton of the handle, also known as core-disc) is the isotropic $i$-disc $0_{D^{i}}\times\{0\}$, while its ascending manifold (i.e. co-core disc) is the coisotropic $2n-i$-disc $D^{*}_{0}D^{i}\times D^{2(n+1-i)}$. Moreover, the Liouville vector-field can be taken to be the radial Liouville vector field corresponding to the Liouville form $-d^{c}\|\mathbf{z}\|^{2}/4$ in the second component.
We begin by symplectically identifying a neighbourhood of the isotropic $k+1$-disc $C_{\eta}$ with a standard Weinstein $k+1$-handle embedded in ×$\mathbb{R}\times Y$, which has been attached along the isotropic sphere $\partial\eta\subset\partial((-\infty,0]\times Y)$. Here the framing of $\partial\eta$ in $Y$ is induced from the framing of $\partial\eta\subset\Lambda$. Then we denote by $\overline{X}_{a}\subset\mathbb{R}\times Y$ the embedded Weinstein sub-cobordism consisting of $(-\infty,0]\times Y$ together with the attached Weinstein $k+1$-handle. This sub-cobordism has a smooth contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ along which the Liouville vector field induced by the handle-attachment is outwards pointing. (Note that this is different from the original Liouville vector field $\partial_{t}$ on the symplectisation.)
Finally, after a compactly supported deformation of the Liouville vector field $\partial_{t}$ of (×$(\mathbb{R}\times Y,d(e^{t}\alpha))$ to a complete Liouville vector field $\zeta$ without critical points for the same symplectic form, we may assume that $L_{\eta}\setminus X_{b}$ is tangent to $\zeta$ in ×$\mathbb{R}\times Y\setminus\overline{X}_{b}$, and that $\zeta$ coincides with the Liouville vector field induced by the Weinstein handle-decomposition of $\overline{X}_{b}$ near its boundary $\partial\overline{X}_{b}$. This gives us the sought Weinstein structure on ×$\mathbb{R}\times Y$ and thus finishes the proof of properties (1)–(4).
  • A Weinstein one-handle attached at the 0-sphere $\partial\eta$;
  • A Weinstein two-handle attached along a knot $S_{\eta}$ that coincides with $\eta$ outside of the above 00-handle; while
In Darboux balls that contain $\partial\eta$ the Legendrian $\Lambda_{\eta}$ and attaching 00-spheres and $1$-spheres $S_{\eta}$ can be described as shown on the bottom in Figure 13.
3.3. Ambient surgery, cusp-connected sum, and isotopies.
Since the main focus here is Legendrian ambient 00-surgery on Legendrian knots in $\mathbb{R}^{3}$ we will give some additional details in this case.
[…]
Figure 12. The Legendrian isotopy of $\Lambda$ following the surgery arc
The isotopy performed in Step (2) is easy to describe also in the front of $\mathbb{R}^{3}$; we express this in a lemma.
Lemma 3.3.
@@
Figure 14. The Kirby diagram on the bottom is obtained from the one on the bottom of Figure 13 by a contact isotopy which is the identity in a neighbourhood of $S_{\eta}$, is supported in the one-handle, and which moves $\Lambda_{\eta}$ so that it enters the one-handle on the same side as $S_{\eta}$, with the two strands having larger $z$–coordinates.
@@
Lemma 4.1.
Proof.
$\Phi_{t}\in\mathrm{Cont}(Y\times T^{*}S^{n}),$
$\Phi_{t}(y,\theta,p)=(\phi_{t}(y),\theta,p\cdot e^{g_{t}(y)})$
@@
Assume that two Legendrian links (resp. union of properly embedded arcs) $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ admit a (compactly supported) Legendrian isotopy $S_{t}\subset Y$, where $\Lambda\subset Y$ is a Legendrian link. It follows that:
@@
    has vanishing algebraic intersection number.
@@
Create Whitney disc by connecting two double points of opposite signs with arcs in both intersecting sheets of the immersed trace cobordism $\mathcal{T}$; this forms a loop in $[0,1]_{t}\times Y\times T^{*}S^{n}$. We may assume that each arc intersects each constant $t$-slice $\{t\}\times Y\times T^{*}S^{n}$ in at most one point. Since the Legendrians have codimension strictly greater than two in the contact manifold $Y\times T^{*}S^{n}$, one can find a smooth family of paths in each slice $\{t\}\times Y\times T^{*}S^{n}$ that connect the two unique points on the arcs in the sheets of $\mathcal{T}$. The high codimension makes it possible to assume that these family of paths trace out an embedded two-dimensional disc with boundary on $\mathcal{T}$ and interior disjoint from $\mathcal{T}$. These Whitney discs can be used to deform $S_{t}\times 0_{S^{n}}$ in the required manner, to remove any intersection with $\Lambda\times 0_{S^{n}}$.
@@
Using the fact that products of stabilised knots become loose [10, Proposition 1.13], combined with Murphy’s $h$-principle for loose Legendrian embeddings [32], we get:
@@
If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.2 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $\Lambda\times 0_{S^{n}}\subset Y_{S_{i}}\times T^{*}S^{n}$, $i=0,1$, are contactomorphic. When $Y_{S_{i}}\cong\mathbb{R}^{3}$, then the induced Legendrians inside $\mathbb{R}^{3}\times T^{*}S^{n}$ are Legendrian isotopic.

### v2

@@ 3.1. Decomposable fillings
A particularly nice class of Lagrangian cobordisms in $\mathbb{R}\times S^{3}$ are the decomposable ones, which were defined by Chantraine in [4]. By definition, a Lagrangian cobordism is decomposable if it consists of a concatenation of the following simple types of Lagrangian cobordisms:
@@
Figure 9. Ambient Legendrian surgery: The Legendrian $\Lambda$ is deformed in a neighborhood of the surgery arc ($\eta$ shown at the bottom) to yield the Legendrian $\Lambda_{+}=\Lambda_{\eta}$ shown on top. There is a standard Lagrangian 1-handle cobordism with concave boundary the Legendrian $\Lambda$ before the surgery, and convex boundary the Legendrian $\Lambda_{+}=\Lambda_{\eta}$ produced by the surgery.
@@
Decomposable Lagrangian cobordisms were shown to be regular in the work of Conway, Etnyre and Tosun [6]. It is still unknown whether the converse is true or not. The proof in [6] moreover gives an explicit recipe for finding the attaching spheres on the complement of the Legendrian. We proceed to explain this recipe in the case of a decomposable Lagrangian filling.
The 0-handles: For each zero-handle we consider $(\mathbb{C}^{2},-d^{c}\rho)$, where $\rho(\mathbf{z})=\frac{\|\mathbf{z}\|^{2}}{4},$ with the standard filling $\mathfrak{Re}\mathbb{C}^{2}\subset\mathbb{C}^{2}$ of the standard Legendrian unknot $\partial_{\infty}\left(\mathfrak{Re}\mathbb{C}^{2}\right)\subset S^{3}$ of $\tt{tb}=-1$. This filling is clearly is tangent to the radial Liouville vector field. In the general case, when the Lagrangian cobordism has several 0-handles, one has to consider a Weinstein connected sum of such Weinstein four-balls. Since this Weinstein connected sum again produces the standard sphere, and since these handles can be determined from the number of zero-handles, we will omit these 1-handles from the diagram.
The 1-handles: We add a pair of cancelling Weinstein one and two-handles just above the region where the cusp-connected sum takes place; this is move K–I shown in Figure 4. The resulting diagram after having performed the cusp-connected sum is shown at the top of Figure 5. Performing the K–II move as shown in the same figure places the two strands of the Legendrian that underwent cusp-connected sum inside the Weinstein 1-handle. (Alternatively, one could handle-slide the two facing cusp-edges together with the surgery arc into the one-handle similarly to the K–II move as shown in Figure 10.) It can readily be seen that the corresponding standard Lagrangian one-handle in this position can be made tangent to the standard Liouville structure on the Weinstein one-handle. In particular, the critical point of the Weinstein one-handle coincides with the critical point of the standard Lagrangian one-handle. See Proposition 3.2 for the general statement in arbitrary dimensions.
@@ 3.2. The case of a general ambient Legendrian surgery
of the zero-section in the above standard neighbourhood, while the Legendrian $\Lambda$ is identified with the Lagrangian co-normal bundle $N^{*}_{(\partial\eta)\times D^{n-k-1}}=N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$ of $(\partial\eta)\times D^{n-k-1}$, lifted to have vanishing $\epsilon,\epsilon]_{z}-$coordinate.
The choice of conformal symplectic normal bundle induces an identification of the $D^{*}D^{n-k-1}$-factor in the above neighbourhood. Since $N^{*}_{(\partial\eta)\times D^{n-k-1}}$ is a trivial $\mathbb{R}$-fibration over $(\partial\eta)\times D^{n-k-1}$, the compatibility of the framing of $\partial\eta$ and the above conformal symplectic normal bundle implies that the $\mathbb{R}$-fibre of the conormal bundle together with the standard coordinates on the $0_{D^{n-k-1}}$-factor together yield the framing of $\partial\eta\subset\Lambda$.
[…]
Passing to the symplectisation, we can find a neighbourhood of $0,1]\times\eta\subset\mathbb{R}\times Y$ in the symplectisation of the form
@@
in which the cone $0,1]\times\eta$ is identified with $0_{[0,1]}\times 0_{\eta}\times\{0\}$ and the cone $0,1]\times\Lambda$ is identified with $0_{[0,1]}\times N^{*}_{\partial\eta}\times 0_{D^{n-k-1}}$. The standard Lagrangian $k+1$-handle is constructed to be a cylinder over $\Lambda$ outside the above neighbourhood of the symplectisation, while the intersection with the above neighbourhood is an explicitly constructed neighbourhood of the core-disc; see [9, Section 4.2.2]. We will only need a description of $L_{\eta}$ in a neighbourhood of the zero-section
@@
Namely, in a small neighbourhood of this zero-section, the standard handle $L_{\eta}$ is given by the Lagrangian co-normal bundle of a $k+1$-dimensional disc
$\overline{C}_{\eta}=\Gamma\times\{0\}\subset[0,1]\times\eta\times D^{n-k-1}$
where $\Gamma\subset[0,1]\times\eta$ is a smooth codimension one embedding of a disc given as the union of $0,1/2]\times\partial\eta$ and a suitable graph of a function $\eta\to[1/2,1]$. In particular, this means that $L_{\eta}$ intersects the zero-section of this standard neighbourhood in a representative $\overline{C}_{\eta}\subset\overline{L}_{\eta}$ of the $k+1$-dimensional core-disc of the handle. Further, we note that the core disc $\overline{C}_{\eta}$ is the boundary of an isotropic $k+2$-disc $\overline{D}_{\eta}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$ given by the sub-level set of the above graph. Here $\overline{D}_{\eta}$ is homeomorphic to a disc and has boundary with corners; one boundary stratum is the disc $\overline{C}_{\eta}$, while the other stratum is $\{0\}\times 0_{\eta}\times\{0\}\subset 0_{[0,1]}\times 0_{\eta}\times\{0\}$. By
@@
    $L_{\eta}$ is contained in the Weinstein $k+1$-handle and is tangent to the Liouville flow, where the critical point of the Liouville flow of the handle coincides with the critical point of the restriction $t|_{L_{\eta}}$ of the symplectisation coordinate;
@@
endowed with a Liouville form for which $\{(0,0)\}$ is a unique critical point of index $i$, whose descending manifold (i.e. the skeleton of the handle, also known as core-disc) is the isotropic $i$-disc $0_{D^{i}}\times\{0\}$, while its ascending manifold (i.e. co-core disc) is the coisotropic 2ni2n-i-disc $D^{*}_{0}D^{i}\times D^{2(n+1-i)}$. Moreover, the Liouville vector-field can be taken to be the radial Liouville vector field corresponding to the Liouville form $d^{c}\|\mathbf{z}\|^{2}/4$ in the second component.
We begin by symplectically identifying a neighbourhood of the isotropic $k+1$-disc $C_{\eta}$ with a standard Weinstein $k+1$-handle embedded in $\mathbb{R}\times Y$, which has been attached along the isotropic sphere $\partial\eta\subset\partial((-\infty,0]\times Y)$. Here the framing of $\partial\eta$ in $Y$ is induced from the framing of $\partial\eta\subset\Lambda$. Then we denote by $\overline{X}_{a}\subset\mathbb{R}\times Y$ the embedded Weinstein sub-cobordism consisting of $(-\infty,0]\times Y$ together with the attached Weinstein $k+1$-handle. This sub-cobordism has a smooth contact boundary $\partial\overline{X}_{a}\subset\mathbb{R}\times Y$ along which the Liouville vector field induced by the handle-attachment is outwards pointing. (Note that this is different from the original Liouville vector field $\partial_{t}$ on the symplectisation.)
@@
Figure 13. The Kirby diagram on the bottom is obtained from the one on the bottom of Figure 12 by a contact isotopy which is the identity in a neighbourhood of $S_{\eta}$, is supported in the one-handle, and which moves $\Lambda_{\eta}$ so that it enters the one-handle on the same side as $S_{\eta}$, with the two strands having larger zz–coordinates.
Finally, after a compactly supported deformation of the Liouville vector field $\partial_{t}$ of $(\mathbb{R}\times Y,d(e^{t}\alpha))$ to a complete Liouville vector field $\zeta$ without critical points for the same symplectic form, we may assume that $L_{\eta}\setminus X_{b}$ is tangent to $\zeta$ in $\mathbb{R}\times Y\setminus\overline{X}_{b}$, and that $\zeta$ coincides with the Liouville vector field induced by the Weinstein handle-decomposition of $\overline{X}_{b}$ near its boundary $\partial\overline{X}_{b}$. This gives us the sought Weinstein structure on $\mathbb{R}\times Y$ and thus finishes the proof of properties (1)–(4).
  • A Weinstein one-handle attached at the 0-sphere $\partial\eta$;
  • A Weinstein two-handle attached along a knot $S_{\eta}$ that coincides with $\eta$ outside of the above 0-handle; while
In Darboux balls that contain $\partial\eta$ the Legendrian $\Lambda_{\eta}$ and attaching 0-spheres and $1$-spheres $S_{\eta}$ can be described as shown on the bottom in Figure 12.
3.3. Ambient surgery, cusp-connected sum, and isotopies.
Since the main focus here is Legendrian ambient 0-surgery on Legendrian knots in $\mathbb{R}^{3}$ we will give some additional details in this case.
The isotopy performed in Step (2) is easy to describe also in the front of $\mathbb{R}^{3}$; we express this in a lemma. The techniques, and moves shown in the figures, are covered by the generalised Reidemeister moves for Legendrian graphs developed in [35], which also have been used by Sabloff, Vela-Vick, Wong, and Wu in [38].
Lemma 3.3.
[…]
Figure 14. The Legendrian isotopy of $\Lambda$ following the surgery arc. These moves were also established by Sabloff, Vela-Vick, Wong and Wu, see e.g. [38, Figure 9] for (S2).
@@
4.1. Generalities on formal Legendrian isotopies in the complement of a Legendrian
@@
Let $\Lambda\subset Y^{3}$ denote a fixed Legendrian link. Further, let $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ be two Legendrians in the complement of $\Lambda$, where each component of $S_{i}$ is either closed or compact with boundary. If there exists a Legendrian isotopy $S_{t}\subset Y$ which takes $S_{0}$ to $S_{1}$ (the isotopy is allowed to cross $\Lambda$), which is the identity near the boundary $\partial S_{0}\cup\partial S_{1}$, then it follows that:
@@
    have vanishing algebraic intersection number.
@@
Create Whitney disc by connecting two double points of opposite signs with arcs in both intersecting sheets of the immersed trace cobordism $\mathcal{T}$; this forms a loop in $0,1]_{t}\times Y\times T^{*}S^{n}$. We may assume that each arc intersects each constant $t$-slice $\{t\}\times Y\times T^{*}S^{n}$ in at most one point. Since the Legendrians have codimension strictly greater than two in the contact manifold $Y\times T^{*}S^{n}$, one can find a smooth family of paths in each slice $\{t\}\times Y\times T^{*}S^{n}$ that connect the two unique points on the arcs in the sheets of $\mathcal{T}$. The high codimension makes it possible to assume that these family of paths trace out an embedded two-dimensional disc with boundary on $\mathcal{T}$ and interior disjoint from $\mathcal{T}$. These Whitney discs can be used to deform $S_{t}\times 0_{S^{n}}$ in the required manner, to remove any intersection with $\Lambda\times 0_{S^{n}}$.
@@
Using the fact that products of stabilised knots become loose [11, Proposition 1.13], combined with Murphy’s $h$-principle for loose Legendrian embeddings [34], we get:
@@
If the two Legendrians $S_{0},S_{1}\subset Y^{3}\setminus\Lambda$ satisfy the assumptions of Proposition 4.1 and, moreover, are stabilised when considered inside the contact three-manifold $Y^{3}\setminus\Lambda$, then the two Legendrians $\Lambda\times 0_{S^{n}}\subset Y_{S_{i}}\times T^{*}S^{n}$, $i=0,1$, are compactly supported contactomorphic. When $Y_{S_{i}}\cong\mathbb{R}^{3}$, then the induced Legendrians inside $\mathbb{R}^{3}\times T^{*}S^{n}$ are Legendrian isotopic.
The existence of a Legendrian isotopy in $\mathbb{R}^{3}\times T^{*}S^{n}$ is a direct consequence of the existence of a compactly supported contact contactomorphism. Namely, since any compact subset in $\mathbb{R}^{3}$ can be displaced from itself by a contactomorphism, any compactly supported contactomorphism of $\mathbb{R}^{3}\times T^{*}M$ preserves the Legendrian isotopy class of any compact Legendrian; see Lemma 4.5 below.